# Confinement-induced collective motion in suspensions of run-and-tumble particles

### v1

@@ II.2.2 Following the structures: Band velocity
The corresponding CM coordinate was then obtained by inverting this transformation back to real space. The coordinate $x_{\mathrm{CM}}(t)$ thus obtained provides a smooth measure of the band’s translational motion, from which the instantaneous and mean propagation velocities were extracted. As shown later, the CM velocity seems to be well defined with a constant value close to $\langle V\rangle\approx\frac{1}{3}\,v_{0}$. We will use this value to characterize the mechanism of the motion.
Figure 2: Phase behavior of the confined active suspension.(a) Dynamic phase diagram in the plane $\phi-l_{p}$ for the active suspension confined in a microchannel with funnel-like obstacles configured at $\theta_{obs}=30^{\circ}$. The different phases are represented by symbols and colors: (b) non-percolated system (blue circles); (c) clogging ( green triangles); and (d) traveling band (red diamonds). The color code of the active particles in b-d represent the orientation angle, $\theta$, with respect to the $x$-axis, according to the color bar on the right. All the graphics of trajectories of simulations presented in this article were generated in part using the visualization software OVITO stukowski2009visualization.
II.2.3 Motion mechanism: Fluxes and internal particle displacements in the travelling band
@@ III.1 Phase behavior
We start by exploring the phase behavior of the active suspension confined in microchannels with $\theta_{obs}=30^{o}$ in the $l_{p}-\phi$ plane (Figure 2a), where the different phases are: non-percolating clustering state (blue circles, panel b), clogging (green triangles, panel c) and traveling band (red diamonds, panel d). Videos of these states are available in the Supplementary Video 1.
@@
To classify, from a structural perspective, the different self-organized steady states observed in the active suspension, we determine the probability of finding a percolating configuration along the $y$-axis at the steady state, $P_{p}$. Thus, high values of $P_{p}$ indicate states where a big dense cluster spans the system in the $y$-direction, whereas small values correspond to states in which the active particles are not forming such structures. The dense structures may correspond to clogs or traveling bands. To distinguish between these two, we monitor the speed of the center of mass of the whole system in the x direction. States corresponding to a clog are characterised by velocities close to zero, in contrast to states corresponding to traveling bands. Combining the percolation and the center of mass velocity criteria, we build the phase diagram in Fig. 2a.
@@
The clustering state consists of active particles accumulating at the funnels or at the flat wall of the microchannel and takes place at low $l_{p}$ and $\phi$ (see Fig. 2b). The clog state mainly occurs for suspensions with high $l_{p}$ and $\phi$ (see Fig. 2c). It also appears when confining the suspension in channels with two flat walls (with the same $l_{p}$ and $\phi$), and can be acknowledged as the confined equivalent of motility-induced phase separation (MIPS) in a bulk suspension with periodic boundary conditions cates2015motility; knippenberg2024motility (See Figure S1a (bulk) and S1b (confinement between flat walls) and Supplementary Video 2). The traveling band is observed in persistent active suspensions ($l_{p}>4\times 10^{3}$) and intermediate to low packing fractions ($0.30\leq\phi\leq 0.20$), and consists of a big cluster of active particles that moves persistently towards left or right. In the example shown in Fig. 2d, it moves to the left.
@@
Interestingly, the traveling band is only stable under confinement in microchannels featuring either funnel-like obstacles with gaps or corrugations. Neither in bulk nor in confinement within channels with two flat walls the band is stable, suggesting that its formation is induced by the asymmetry of the obstacles in one of the confining walls See Figure S1 and Supplementary Video 2. We also observe that, at low packing fractions, ($\phi\leq 0.25$), the system exhibits a clustering state for low $l_{p}$ that directly becomes a traveling band as $l_{p}$ increases. Contrastingly, at higher packing fractions, ($\phi\geq 0.275$), the system shows clogging behavior before the traveling band is formed as $l_{p}$ increases (See Fig 2a).
@@ III.2 Characterization of the traveling band
Having explored the phase behavior of the system, we now focus on the features of the traveling band. First, we calculate the speed of the center of mass in the $x$ direction of the system, $V_{CM}$, for cases in which a clog or a traveling band is formed. Note that the center of mass of the system roughly matches that of the largest cluster when the system is in a clogging or a traveling state.
@@
In Fig. 3a, we present the probability distributions of $V_{CM}$. The distributions calculated from simulations (circles) fit well to Gaussian distributions (solid lines). The color bar represents the persistence length, $l_{p}$. For clogging states, $P(V_{CM})$ is centered at 0 with a small standard deviation, whereas, for traveling states, the distributions are shifted to non-zero values with higher standard deviations. Clearly, the mean speed of the center of mass, $\langle V_{CM}\rangle$, can be used as an order parameter to monitor the transition from clogging to traveling states, as presented in Fig. 3b. Remarkably, the traveling band reaches average speeds, $\langle V_{CM}\rangle$, of up to $35\%$ as high as the active particles’ self-propulsion speed (See Fig. 3b).
@@
Figure 4: Probability distributions of the orientation angle , $\theta$, of active particles for four selected cases. (a) Clogging states. Filled markers correspond to $\phi=0.40$, and empty to $\phi=0.35$. (b) Traveling states. Filled markers correspond to $\phi=0.275$ and empty to $\phi=0.25$. (c) Orientation domains within the band. I: Particles pointing mainly upwards. II: Particles pointing towards the direction of motion of the traveling band, constituting the thrust. III: Particles forming a thin layer in the lower part of the band pointing opposite to the band’s motion. This domain is the facilitator of the collective motion. IV: Particles trapped in the funnels mainly pointing downwards. The distributions are computed using only the orientations of the largest cluster in the system. The band moves to the right and $\theta_{obs}=30^{\circ}$.
In Fig. 4a we report the $P(\theta)$ for clogging states at different $l_{p}$ and $\phi$, observing, as could be expected, that the distributions are rather uniform. This is already visible from the snapshot presented in Fig. 2c, where, at the clog boundaries, active particles point in opposite directions, while inside the clog they are oriented at random. The profile of the clog boundaries seems parabolic, which could be due to the effective attraction towards the walls induced by activity. On the other hand, the $P(\theta)$ for traveling states in Fig. 4b, also at different $l_{p}$ and $\phi$, exhibits two distinctive peaks around $\theta=90^{o}$ and $\theta=180^{o}$. We have chosen cases in which the traveling structure moves to the right as shown in Fig. 4c, although the simulations show no clear preference for any direction. As seen in Fig. 4b, despite the different values of $l_{p}$ and $\phi$, as long as those values are within the stability region of the traveling band, the distribution of orientations inside the band remains the same, indicating the same underlying mechanism.
@@
III.3 Mechanism of travelling band’s motion
@@
As domain I is the thrust of the band, the counter-flow of particles represented in domainIII is the facilitator of its motion. If particles in domain III could not flow through the lower part of the band, it would lead to the accumulation of particles opposing to the band’s motion in its frontal part, ending up in a clogging state. Particles in domain III are able to find their way out of the band due to the presence of the obstacles that comprise two stable positions for active particles with a height difference (convex and concave gaps), which facilitate hopping and sliding. This is confirmed by simulations of confined active suspensions inside channels with two flat walls, where only clogging states form (See Figure S1).
Figure 5: (a) Representative particle trajectories depending on their initial positions along the $y$-axis inside the traveling band, for $l_{p}=10^{5}$, $\phi=0.3$ and $\theta_{obs}=30^{\circ}$. The wall particles are not shown for the sake of clarity. (b) Corresponding local density profiles $\rho(y)$, total flux $j_{x}(y)$, and counter-current $\Delta(y)$ along the $y$-axis. The channel walls are depicted in yellow for reference.
Since the motion of the bands implies two currents of particles moving in opposite directions- domains I and II in the band’s direction and III in the opposite direction - it somehow resembles the tracked locomotion performed by some vehicles, although both currents move at different speeds (as will be discussed later in Fig. II.2.3). Thus, we name this new type of collective motion confinement-induced tracked locomotion.
Active systems are well known to exhibit strong finite size effects, especially those forming traveling structures kursten2020dry; chate2020dry; fernandez2025dynamics. With the intention of testing the robustness of our results, we perform simulations setting the conditions we expect a traveling band, but for channels with $l_{p}<L_{x}<10l_{p}$. We observe that the traveling bands are still formed and move following the same mechanism described above. However, the traveling bands are stable in a shorter length scale and thus more difficult to analyze See Fig. S3. We are interested in characterizing and unraveling the mechanism of confinement-induced tracked locomotion, so we set the simulations at conditions at which we can observe it for longer time scales ($L_{x}\in[790,1000]$) to have statistically significant results.
@@
In Figure 5b, we present the three averaged quantities involved in equation 5, for a typical system exhibiting confinement-induced tracked locomotion: local density profile, $\rho(y)$ (in black); flux of particles in the $x$ direction along the $y$-axis, $j_{x}(y)$ (in blue); and the counter-current of particles, $\Delta(y)$ (in red). The funnel-like obstacles and the top wall of the channel are depicted in the figure in yellow, for reference. In the concave regions of the obstacles, $\rho(y)$ adopts values close to zero. This is because only one particle per $l_{obs}$ is counted due to the shape of the obstacles. Then, $\rho(y)$ grows fast, up to positions just above the convex gaps of the obstacles (spanning domains IV and II), and then it grows linearly, with a smaller slope, up to the top wall of the microchannel (domains I and II).
As expected, $j_{x}(y)$ shows a small but negative value in the region that corresponds to the flow of particles belonging to domain III in the opposite direction of the band’s motion, whereas the region comprising domains I and II ($y\geq 7.5$) shows a constant high flux of particles in the same direction of the band’s motion. Both the tendencies of $\rho(y)$ and $j_{x}(y)$ are consistent with the structure and dynamics of the band as shown in Fig. 5a (see also Supplementary Video 2).
@@
Finally, $\Delta_{x}(y)$ shows a peak centered at the region corresponding to domain III, whereas it vanishes in the rest of the microchannel. As shown in a previous work fernandez2025dynamics, a non-zero counter-current term, $\Delta_{x}(y)$, signals a source-sink effect that, in the case of confinement-induced tracked locomotion, occurs in the lower part of the band and reflects the dynamics of domain III: particles are captured from at front end of the band and expelled from the back end of the band. Additionally, $\Delta_{x}(y)$ contains all the contributions that oppose to the advective transport represented by the collective motion of the band ($\langle V_{CM}\rangle\neq 0$), thus accounting for the vacancy diffusion of particles due to rearrangements within the band and facilitated by sliding along the obstacles. The vacancy diffusion mechanism that contributes to the motion of particles in domain III is justified since the particles inside the band are arranged in a hexagonal lattice that is continuously distorted, generating empty lattice spaces that are subsequently occupied by other particles.
III.4 Stability of the band with respect to $\theta_{obs}$
With the goal of understanding the role of $\theta_{obs}$ in confinement-induced tracked locomotion, we perform simulations at $\phi=0.30$ and $l_{p}=10^{5}$ for $\theta_{obs}\in[0^{o},75^{o}]$. In Fig. 6, using $\langle V_{CM}\rangle$ as order parameter, we observe the existence of a range $15^{o}\leq\theta_{obs}\leq 45^{o}$ which allows the formation of traveling structures.
Figure 6: Center-of-mass velocity as a function of corrugation tilt angle. Inset represents the dependence of the x-component of the counter-current y-axis $\Delta_{x}(y)$ on tilt angle, when the traveling band appears.

### v2

@@ II.2.2 Following the structures: Band’s velocity
The corresponding CM coordinate was then obtained by inverting this transformation back to real space. The coordinate $x_{\mathrm{CM}}(t)$ thus obtained provides a smooth measure of the band’s translational motion, from which the instantaneous and mean propagation velocities were extracted. As shown later, the CM velocity seems to be well defined with a constant value close to $\langle V\rangle\approx\frac{1}{3}\,v_{0}$. We will use this value to characterize the motion’s mechanism.
Figure 2: Phase behavior of the confined active suspension.(a) Dynamic phase diagram in the plane $\phi-l_{p}$ for the active suspension confined in a microchannel with funnel-like obstacles configured at $\theta_{obs}=30^{\circ}$ and $L_{x}=797\sigma_{22}$. The different phases are represented by symbols and colors. The areas corresponding to each phase have been colored as a rough estimate to provide a guide for the eye. (b) Non-percolated system (blue circles). (c) Clogging ( green triangles). (d) Traveling band (red diamonds). The color code of the active particles in b-d represents the active particle’s orientation angle, $\theta$, with respect to the $x$-axis, according to the color bar on the right. (e) Evolution of the $x$ component of the center of mass of the system, $x_{cm}$, at $l_{p}=10^{5}$ and different packing fractions, $\phi$. (f) Average center of mass velocity of the system, $\langle V_{CM}\rangle$, at persistent length $l_{p}=10^{4}$ and packing fraction $\phi=0.3$ for different channel lengths. Green stars markers used for long channels represent systems where more than two traveling bands form, very often moving in opposite directions. At long times, these traveling bands merge to form clogs. (g) Average percolation probability, $P_{p}$, as a function of $l_{p}$ at $\phi=0.3$. The color of the points correspond to the phases presented in panel a.
II.2.3 Motion mechanism: Fluxes and internal particle displacements in the traveling band
@@ III.1 Phase behavior
We start by exploring the phase behavior of the active suspension confined in microchannels with $\theta_{obs}=30^{o}$ in the $l_{p}-\phi$ plane ($l_{p}=\frac{F_{a}}{\alpha_{r}\gamma}$, $\phi=\frac{N\pi\sigma_{11}^{2}}{4L_{x}L_{y}}$) , as presented in Figure 2a, where the different phases are: non-percolating clustering state (blue circles, panel b), clogging (green triangles, panel c) and traveling band (red diamonds, panel d). To construct the phase diagram, we consider channels with $L_{x}=797\sigma_{22}$. Videos of these states are available in the Supplementary Video 1. All the graphics of trajectories of simulations presented in this article were generated in part using the visualization software OVITO Stukowski (2009).
@@
To classify, from a structural perspective, the different self-organized steady states observed in the active suspension, we determine the probability of finding a percolating configuration along the $y$-axis at the steady state, $P_{p}$. High values of $P_{p}$ indicate states where a big dense cluster spans the system in the $y$-direction, whereas low values correspond to states in which the active particles are not forming such structures. In Figure 2g, we present $P_{p}$ as a function of $l_{p}$ for $\phi=0.30$, illustrating the different self-organized states with the same color code as in the phase diagram (see Fig. 2a) . The dense structures ($P_{p}\approx 1.0$) may correspond to clogs or traveling bands. To distinguish between these two, we monitor the speed of the center of mass of the whole system in the x direction. Figure 2e shows the temporal evolution of the $x$ component of the position of the center of mass, $x_{cm}$. States corresponding to a clog are characterized by constant $x_{cm}$ and thus by velocities close to zero, in contrast to states corresponding to traveling bands , in which $x_{cm}$ exhibits well-defined slopes and thus velocities different than zero. Combining the percolation and the center of mass velocity criteria, we build the phase diagram in Fig. 2a. Note that the traveling states presented in Fig. 2e (in red), correspond to bands moving to the right. However, we observe that the traveling bands move indistinctly to the right and left, as the microchannel geometry lacks features that would bias their motion. We will discuss this in more detail later.
@@
The clustering state consists of active particles accumulating at the funnels or at the flat wall of the microchannel and takes place at low $l_{p}$ and $\phi$ (see Fig. 2b). The clog state occurs for suspensions with high $\phi$ (see Fig. 2c). As $l_{p}$ increases, the formation of the traveling structure is favored, especially for intermediate packing fractions ($\phi\leq 0.325$). The clogging state also appears when confining the suspension in channels with two flat walls (with the same $l_{p}$ and $\phi$), and can be acknowledged as the confined equivalent of motility-induced phase separation (MIPS) in a bulk suspension with periodic boundary conditions Cates and Tailleur (2015); Knippenberg et al. (2024) (see Figure S1a, Figure S1b and Supplementary Video 2). The traveling band is observed in persistent active suspensions ($l_{p}>4\times 10^{3}$) and intermediate to low packing fractions ($0.30\leq\phi\leq 0.20$), and consists of a big cluster of active particles that moves persistently towards left or right. In the example shown in Fig. 2d, it moves to the left.
@@
Interestingly, the traveling band is only stable under confinement in microchannels featuring either funnel-like obstacles with gaps or corrugations. Neither in bulk nor in confinement within channels with two flat walls the band is stable, suggesting that its formation is induced by the asymmetry of the obstacles in one of the confining walls (see Figure S1 and Supplementary Video 2). We also observe that, at low packing fractions, ($\phi\leq 0.25$), the system exhibits a clustering state for low $l_{p}$ that directly becomes a traveling band as $l_{p}$ increases. Contrastingly, at higher packing fractions, ($\phi\geq 0.275$), the system shows clogging behavior before the traveling band is formed as $l_{p}$ increases (see Fig 2a).
Finally, we study the stability of the traveling band as a function of the channel’s length, $L_{x}$. In Figure 2f, we illustrate the average speed of the center of mass of the system, $\langle V_{CM}\rangle$, upon increasing $L_{x}$, for $\phi=0.30$ and $l_{p}=10^{4}$, conditions at which the traveling band is stable as shown in Fig. 2a. For short channels ($L_{x}\lessapprox 500\sigma_{22}$), the traveling structure does not fully develop. In the range $500\sigma_{22}\lessapprox L_{x}\lessapprox 2000\sigma_{22}$, one traveling structure fully develops and exhibits persistent motion for long time scales. For $L_{x}\gtrapprox 5000\sigma_{22}$ we observe $\langle V_{CM}\rangle$ close to zero. However, these states do not strictly correspond to clogs, since many traveling structures form at the beginning of the simulations, moving, very often, in opposite directions. As time evolves, these traveling structures eventually merge into clogs. And thus $V_{CM}$ approaches zero. We mark these states in Fig.2f with green, the characteristic color of clogging states in the phase diagram of Fig. 2a, but with stars instead of triangles, indicating that at the beginning of the simulations the traveling structures are formed with a mean lifetime of a few thousands of $\tau$. Although these traveling bands observed in long channels survived for shorter time scales, their mechanism of motion is exactly the same as the one exhibited by the bands formed in shorter channels with $500\sigma_{22}\lessapprox L_{x}\lessapprox 2000\sigma_{22}$ in which only one stable band is formed. Since the objective of this article is to characterize the mechanism of this novel collective motion induced by confinement, we keep $L_{x}$ within the range in which only one stable traveling band is observed, and thus have, in turn, enough statistics for the analysis.
@@ III.2 Characterization of the traveling band
Having explored the phase behavior of the system, we now focus on the features of the traveling band. First, we calculate the speed of the system’s center of mass in the $x$ direction, $V_{CM}$, for cases in which a clog or a traveling band is formed. Note that the center of mass of the system roughly matches that of the largest cluster when the system is in a clogging or a traveling state.
@@
In Fig. 3a, we present the probability distributions of $V_{CM}$. The distributions calculated from simulations (circles) fit well to Gaussian distributions (solid lines). The color bar represents the persistence length, $l_{p}$. For clogging states, $P(V_{CM})$ is centered at 0 with a small standard deviation, whereas, for traveling states, the distributions are shifted to non-zero values with higher standard deviations. Clearly, the mean speed of the center of mass, $\langle V_{CM}\rangle$, can be used as an order parameter to monitor the transition from clogging to traveling states, as presented in Fig. 3b. Remarkably, the traveling band reaches average speeds, $\langle V_{CM}\rangle$, of up to $35\%$ as high as the active particles’ self-propulsion speed (see Fig. 3b). It should be noted that this motion in one direction has been studied without obstacles in systems with active-passive mixtures Fernandez-Quevedo et al. (2025) or in the presence of activity with an asymmetric spatial dependence in the system Metzger et al. (2024, 2025), such that a traveling band is observed. In the latter cases, a preferred direction emerges in the system, leading to a ratchet effect on the motion.
In contrast to the literature, the system studied here does not exhibit left-right asymmetry, and the appearance of a net motion was not obviously expected. The motion is not induced by a ratchet effect, since at the limit of $t\to\infty$ we would expect zero flux, as shown by direction reversals that compensate for long enough time scales (see Fig S4). However, we stress that the observed traveling states can be very persistent, with a time life of hundreds of thousands of simulation time units, as shown in Fig. S4. If we consider that the moving band typically has a velocity of $\langle V_{CM}\rangle\approx v_{0}/3=\sigma_{11}/(6\,\tau)$ and can persist in a given direction for $\sim 150,000\,\tau$ (see Fig S4), then the entire band can travel $\sim 22,500$ body lengths of active particles, which, in terms of any microorganism performing a run-and-tumble motion, represents an enormous collective persistence.
@@
Figure 4: Probability distributions of the orientation angle , $\theta$, of active particles for four selected cases. The scheme of the active particle shows the definition of $\theta$ as the angle between the orientation vector of the active particle with respect to the x-axis. (a) Clogging states. Filled markers correspond to $\phi=0.40$, and empty to $\phi=0.35$. (b) Traveling states. Filled markers correspond to $\phi=0.275$ and empty to $\phi=0.25$. (c) Orientation domains within the band. I: Particles pointing mainly upwards , acting as a break of the traveling band. II: Particles pointing towards the direction of motion of the traveling band, constituting the thrust. III: Particles forming a thin layer in the lower part of the band pointing opposite to the band’s motion. This domain is the facilitator of the collective motion. IV: Particles trapped in the funnels mainly pointing downwards. The distributions are computed using only the orientations of the largest cluster in the system. The band moves to the right and $\theta_{obs}=30^{\circ}$.
In Fig. 4a we report $P(\theta)$ for clogging states at different $l_{p}$ and $\phi$, observing, as could be expected, that the distributions are rather uniform. This is already visible from the snapshot presented in Fig. 2c, where, at the clog boundaries, active particles point in opposite directions, while inside the clog they are oriented at random. The profile of the clog boundaries seems parabolic, which could be due to the effective attraction towards the walls induced by activity. On the other hand, the $P(\theta)$ for traveling states in Fig. 4b, also at different $l_{p}$ and $\phi$, exhibits two distinctive peaks around $\theta=90^{o}$ and $\theta=180^{o}$. We have chosen cases in which the traveling structure moves to the right as shown in Fig. 4c, although the simulations show no clear preference for any direction. As seen in Fig. 4b, despite the different values of $l_{p}$ and $\phi$, as long as those values are within the stability region of the traveling band, the distribution of orientations inside the band remains the same, indicating the same underlying mechanism.
@@
III.3 Mechanism of traveling band’s motion
@@
As domain II is the thrust of the band, the counter-flow of particles represented in domain III is the facilitator of its motion. If particles in domain III could not flow through the lower part of the band, usually because of an internal rearrangement of the band’s structure, it would lead to the accumulation of particles opposing to the band’s motion in its frontal part, balancing out the thrust force exerted by domain II and, in consequence, ending up in a clogging state. A detailed study on the dynamics of the traveling-clogging transition is left for future work. Particles in domain III are able to find their way out of the band due to the presence of the obstacles that comprise two stable positions for active particles with a height difference (convex and concave gaps), which facilitate hopping and sliding. This is confirmed by simulations of confined active suspensions inside channels with two flat walls, where only clogging states form (see Figure S1).
Since the motion of the bands implies two currents of particles moving in opposite directions- domains I and II in the band’s direction and III in the opposite direction - it somehow resembles the tracked locomotion performed by some vehicles, although both currents move at different speeds (as will be discussed later in Fig. 5). Thus, we name this new type of collective motion confinement-induced tracked locomotion.
Active systems are well known to exhibit strong finite size effects, especially those forming traveling structures Kürsten and Ihle (2020); Chaté (2020); Fernandez-Quevedo et al. (2025). With the intention of testing the robustness of our results, we perform simulations setting the conditions we expect a traveling band, but for channels with $l_{p}<L_{x}<10l_{p}$. We observe that the traveling bands are still formed and move following the same mechanism described above. However, the traveling bands are stable in a shorter time scale and thus more difficult to analyze (see Fig. S3). The reason behind this reduced stability is the formation of many traveling bands within the long microchannel, some of them moving in opposite directions, that eventually collide to form clogs.
Figure 5: (a) Representative particle trajectories depending on their initial positions along the $y$-axis inside the traveling band, for $l_{p}=10^{5}$, $\phi=0.3$ and $\theta_{obs}=30^{\circ}$. Wall particles are not shown for the sake of clarity. (b) Corresponding local density profiles $\rho(y)$, total flux $j_{x}(y)$, and counter-current $\Delta_{x}(y)$ along the $y$-axis. The channel walls are depicted in yellow for reference.
Figure 6: Center-of-mass velocity as a function of: (a) Corrugation tilt angle when $L_{y}=10\sigma_{11}$ and $\phi=0.30$. Inset represents the dependence of the x-component of the counter-current along the y-axis, $\Delta_{x}(y)$, on tilt angle, when the traveling band is formed; (b) top wall’s corrugation angle, $\beta$ (see Supplementary Video 4); and (c) the width of the channel, $L_{y}$, when $\theta_{obs}=30^{o}$ for $\phi=0.30$ and $0.50$. The lines are a guide for the eyes to correctly observe trends. Note that the angle $\beta$ defines a gradual transformation of the top wall from flat to a trapezoidal wave which is half a period out of phase with respect to the funnel-like obstacles in the bottom wall, thus the maxima of the top wall coincide with the minima of the bottom wall and vice versa.
@@
In Figure 5b, we present the three averaged quantities involved in equation 5, for a typical system exhibiting confinement-induced tracked locomotion: local density profile, $\rho(y)$ (in black); flux of particles in the $x$ direction along the $y$-axis, $j_{x}(y)$ (in blue); and the counter-current of particles, $\Delta_{x}(y)$ (in red). The funnel-like obstacles and the top wall of the channel are depicted in the figure in yellow, for reference. In the concave regions of the obstacles, $\rho(y)$ adopts values close to zero. This is because only one particle per $l_{obs}$ is counted due to the shape of the obstacles. Then, $\rho(y)$ grows fast, up to positions just above the convex gaps of the obstacles (spanning domains IV and II), and then it grows linearly, with a smaller slope, up to the top wall of the microchannel (domains I and II). As expected, $j_{x}(y)$ shows a small but negative value in the region that corresponds to the flow of particles belonging to domain III in the opposite direction of the band’s motion, whereas the region comprising domains I and II ($y\geq 3.75$) shows a constant high flux of particles in the same direction of the band’s motion. Both the tendencies of $\rho(y)$ and $j_{x}(y)$ are consistent with the structure and dynamics of the band as shown in Fig. 5a (see also Supplementary Video 3).
@@
Finally, $\Delta_{x}(y)$ shows a peak centered at the region corresponding to domain III, whereas it vanishes in the rest of the microchannel. As shown in a previous work Fernandez-Quevedo et al. (2025), a non-zero counter-current term, $\Delta_{x}(y)$, signals a source-sink effect that, in the case of confinement-induced tracked locomotion, occurs in the lower part of the band and reflects the dynamics of domain III: particles are captured at the front end of the band and expelled from the back end of the band. Additionally, $\Delta_{x}(y)$ contains all the contributions that oppose to the advective transport represented by the collective motion of the band ($\langle V_{CM}\rangle\neq 0$), thus accounting for the vacancy diffusion of particles due to rearrangements within the band and facilitated by sliding along the obstacles. The vacancy diffusion mechanism that contributes to the motion of particles in domain III is justified since the particles inside the band are arranged in a hexagonal lattice that is continuously distorted, generating empty lattice spaces that are subsequently occupied by other particles.
III.4 Stability of the band with respect to confining geometry
With the goal of understanding the role of confinement in the emergence of traveling states, we now focus on varying the geometrical parameters of the confining microchannels. In particular, we study the stability of the traveling structure when changing $\theta_{obs}$ while keeping the top wall flat (see Fig. 6a); changing the angle $\beta$ that controls the shape of trapezoids in the top wall as shown in the inset of Fig. 6b while $\theta_{obs}=30^{o}$ is maintained constant; and changing the width of the channel $L_{y}$ whereas $\theta_{obs}=30^{o}$ is kept constant and the top wall kept flat.